# Understanding the Ranking Loss for Recommendation with Sparse User Feedback

### v1

@@ 4.5. RQ4: Beyond Ranking Loss
(a) Focal Loss
(b) Negative Sampling
Figure 7. Evaluation of Focal Loss and negative sampling. (a) Comparison of Gradient Norm (boxplot) and AUC (solid line) under different $\gamma$ for Focal loss with BCE loss. (b) AUC (blue) and Logloss (red) with increasingly aggressive negative sampling.
@@
We’re curious whether other approaches beyond ranking loss can alleviate gradient vanishing and, therefore, improve performance by alleviating the gradient vanishing of negative samples. We pick up Focal Loss (Lin et al., 2017), which assigns higher weights to poorly-classified samples, i.e., negative samples suffering from gradient vanishing in our scenario. Specifically, Focal Loss introduces a weight with hyper-parameter $\gamma$ to control the weight of samples:
(15) $\mathcal{L}_{Focal}=-\frac{1}{N}\sum_{i\in N}[y_{i}(1-\hat{p}_{i})^{\gamma}\log(\hat{p}_{i})+(1-y_{i})\hat{p}_{i}^{\gamma}\log(1-\hat{p}_{i}))],$
@@
where $\gamma$ controls the relative weight. For those negative samples that suffered from gradient vanishing, their prediction score $\hat{p}_{i}$ should be wrongly high, making them have a higher weights $\hat{p}_{i}^{\gamma}$ than those negatives that are well-classified, i.e., with a low score. The larger $\gamma$, the higher the relative weights to those poorly-classified samples.
@@
We then conducted a set of experiments by comparing gradients and performance for Focal Loss with different $\gamma$. As shown in Fig. 7(a), similar to Combined-Pair, Focal Loss also gets higher gradients on negative samples than the BCE loss. The larger the $\gamma$, the more weights on the poorly classified samples, and the larger the performance lift than the BCE method. These phenomena further validate our perspective that the gradient vanishing of negative samples is the key challenge when positives are sparse; therefore, resolving it with either ranking loss or Focal Loss leads to performance lift.
@@
Please note that the original $\hat{p_{i}}^{\gamma}$ is always less than $1.0$, and the gradients of negative samples in Focal loss are constantly smaller than BCE loss. For a fair comparison, we introduce a slight modification to the original formulation by replacing $\hat{p_{i}}^{\gamma}$ with $\alpha(\hat{p_{i}})$. This adjustment normalizes the average weights of the negative samples to $1.0$, aligning it with the BCE loss. Formally, the modified version is defined as:
(16) $\mathcal{L}_{Focal^{\prime}}=-\frac{1}{N}$ $\sum\limits_{i\in N}[y_{i}(1-\hat{p}_{i})^{\gamma}\log(\hat{p}_{i})$
$+$ $[\hat{p_{i}}^{\gamma}+(1-\sum_{k\in N_{-}}\hat{p_{k}}^{\gamma}/N_{-})](1-y_{i})\log(1-\hat{p}_{i}))],$
@@
Another trial is to reduce the proportion of negative samples through negative sampling, thereby increasing the estimated CTR. This may consequently increase the gradient of negative samples because their gradients are proportional to the estimated CTR (Eq. 3.2.1). However, optimizing quality is not only related to gradient issues but also to the quality of negative samples. Some studies (Chen et al., 2020) have shown that negative samples of appropriate scale and learning difficulty contribute to learning representations of uniformity. Additionally, discarding negative samples can lead to information degradation, especially in datasets like Criteo that have already undergone negative sampling. Hence, we observe that as negative sampling becomes more aggressive, both the model’s ranking and classification capabilities deteriorate in Fig. 7(b), which indicates that negative sampling fails to improve performance.
@@ 5. Stability and Compatibility
Recent studies(Bai et al., 2023; Yan et al., 2022) have raised concerns regarding tailoring ranking loss into CTR prediction. Some studies (Yan et al., 2022) are concerned that ranking loss introduces score drifting, leading to optimization instability, while others (Bai et al., 2023) express concerns about compatibility issues between ranking loss and BCE loss. Both issues can also lead to negative effects on optimization. Hence, we aim to investigate whether classification-ranking loss, especially Combined-Pair, also has these two issues.
@@ 5.1. Stablility
Stability (Yan et al., 2022) refers to the phenomenon that the model scores may keep drifting during model training. Intuitively, singly employing ranking loss causes the model to focus solely on the relative orders between samples, neglecting the absolute prediction scores. This leads to score drifting (Yan et al., 2022), resulting in non-scale calibrated outcomes and exacerbating bias distribution. Hence, we wonder whether combining ranking loss with BCE loss in Combined-Pair may worsen the bias compared to using only BCE loss (i.e., the BCE method). It is worth mentioning that we did not employ any post-processing calibration techniques; otherwise, the bias would not reflect its training stability.
@@
We bucketize the samples with equal frequency of positive samples and plot the bias of corresponding samples within each bucket. As shown in Fig. 8, in both online and offline advertising, the BCE method severely underestimates the click-through rate (CTR) in the lower buckets, while Combined-Pair has a much smaller bias. In the higher buckets, the BCE method overestimates the CTR, and Combined-Pair also has a small bias in those buckets. The reason is that Combined-Pair mitigates the gradient vanishing issue, thus facilitating model optimization and finally leading to lower bias. In summary, Combined-Pair demonstrates superior calibration ability compared to the BCE method, avoiding the issue of score drifting and ensuring stability.
@@ 5.2. Compatibility
Recent research (Bai et al., 2023) argued that the pointwise and pairwise loss may not be compatible with each other since they have different global minima. Thus, we are curious about the compatibility of the two losses within Combined-Pair: the RankNet loss v.s. the BCE loss. We refrain from comparing their global minima (Bai et al., 2023), as differences in global minima among losses do not inherently indicate incompatibility. For instance, while l1 or l2 regularization losses do not share the same global minimum as most optimization objectives, they are still effectively utilized in tandem to mitigate overfitting.
Instead, we investigate the compatibility of these losses by analyzing their gradients, examining whether there is a conflict in gradients between them. As shown in Tab. 2, the gradients of BCE loss and RankNet loss are always aligned in the same direction. Therefore, the optimization directions of the combined-pair loss remain consistent.
@@
Negative $\frac{\hat{p}_{j}}{N}$ $\frac{\sum\limits_{i\in N_{+}}\sigma(z_{j}^{(-)}-z_{i}^{(+)})}{N_{+}N_{-}}$ Same ($>0$)
Positive $-\frac{1-\hat{p}_{i}}{N}$ $-\frac{\sum\limits_{j\in N_{-}}[1-\sigma(z_{i}^{(+)}-z_{j}^{(-)})]}{N_{+}N_{-}}$ Same ($<0$)
@@ 6. Online Deployment
We conduct an empirical evaluation of Combined-Pair on the Click-Through Rate (CTR) prediction task across three distinct online advertising scenarios in Tencent: WeChat Channels, WeChat Moments, and the Demand-Side Platform (DSP). The underlying model architecture utilizes Heterogeneous Experts with Multi-Embedding framework (Guo et al., 2023b; Su et al., 2023; Pan et al., 2024). Specifically, our approach involves training multiple feature interaction experts, such as GwPFM (Pan et al., 2024) (a variant of FFM (Juan et al., 2017) and FwFM (Pan et al., 2018)), IPNN (Qu et al., 2016), DCN V2 (Wang et al., 2021), or FlatDNN, to capture diverse feature interactions for sparse ID features. Additionally, we learn multiple embedding tables for all features, with each table corresponding to one or several experts. For sequence features, we employ the TIN (Zhou et al., 2023) model, which leverages target attention mechanisms (Zhou et al., 2018, 2019) to learn the semantic-temporal correlation.
@@
Based on the above backbone model architecture, we deployed Combined-Pair and conducted online A/B testing from early July 2023 to August 2023. We configured the RankNet Loss with a weight of 0.1 and adopted streaming training. The CTR varies from $0.1\%$ to $2.0\%$ in different scenarios and shows extreme sparsity. We examined the gradient distribution of negative samples and normalized their frequency as shown in Fig. 9. Our findings indicate that the gradient distribution of negative samples for the Combined-Pair method is significantly right-skewed compared to the BCE method, demonstrating that the Combined-Pair mitigates gradient vanishing in online advertising.
Upon further examination of the model’s online performance metrics, we observed that Combined-Pair significantly improves the effectiveness compared to the BCE method, as shown in Tab. 3. For example, during one month’s A/B testing with 20% traffic, Combined-Pair achieves $0.59\%$ cost lift and $0.70\%$ GMV (Gross Merchandise Value) lift over the BCE method.
Besides, for new ads, few training samples and, thus, the lack of interactions cause insufficient gradient updates and more severe gradient vanishing problems. Therefore, we pay attention to the new ads and find that Combined-Pair obtains a significant performance improvement in the WeChat Channels scenario, with a $1.26\%$ GMV lift and $0.34\%$ cost lift, as shown in Tab. 4, The improvement is statistically significant according to the t-test. We estimate this performance gain leads to hundreds of millions of dollars in revenue lift per year. Combined-Pair has been successfully deployed in our production system in September 2023.
Based on the aforementioned backbone model architecture, we deployed Combined-Pair and conducted online A/B testing from early July 2023 to August 2023. During this period, we configured the RankNet Loss with a weight of 0.1 and employed streaming training. It is worth noting that the Click-Through Rate (CTR) exhibited significant variation, ranging from 0.1% to 2.0%, across different scenarios, indicating extreme sparsity.
To gain insights into the gradient distribution of negative samples, we examined and normalized their frequency, as depicted in Figure 1. Our analysis revealed that the gradient distribution of negative samples for the Combined-Pair method exhibited a pronounced right-skewness compared to the BCE method. This observation provides compelling evidence that Combined-Pair effectively mitigates the issue of gradient vanishing in the context of online advertising.
Further evaluation of the model’s online performance metrics demonstrated the significant effectiveness improvement achieved by Combined-Pair compared to the BCE method, as summarized in Table 1. For instance, during one month of A/B testing with 20% traffic allocation, Combined-Pair yielded a notable cost lift of 0.59% and a Gross Merchandise Value (GMV) lift of 0.70% over the BCE method.
Moreover, we paid particular attention to the performance of Combined-Pair for new ads, which often suffer from insufficient training samples and limited interactions, leading to more severe gradient vanishing issues. Our analysis revealed a substantial performance improvement in the WeChat Channels scenario, with a GMV lift of 1.26% and a cost lift of 0.34%, as presented in Table 2. Importantly, statistical analysis using the t-test confirmed the significance of this improvement. We estimate that this performance gain translates to a revenue lift of hundreds of millions of dollars per year.
@@
Following the discussion of the classification ability of Combined-Pair, we want to examine the LogLoss of online deployment. We observed that the reduction of LogLoss of Combined-Pair ranges from 0.01% to 0.1% in 7 days compared with the BCE method during the A/B testing period, which validates the previous discussion that Combined-Pair could improve the model’s classification ability.
@@ 7. Conclusion
In online advertising, incorporating ranking loss upon the classification loss has demonstrated superior performance. In this paper, we discover a challenge of only using classification loss for recommendation when the positive feedback is sparse: the gradient vanishing of negative samples. We then introduce a novel perspective to explain the effectiveness of the classification-ranking loss: the incorporated ranking loss introduces large gradients for negative samples, hence mitigating their gradient vanishing issue. Comprehensive experiments and analysis provided strong evidence in support of our perspective. We hope that this gradient-based perspective can contribute to the optimization of CTR prediction models and inspire further research in this field.

### v2

@@ 4.5. RQ4: Beyond Ranking Loss
We’re curious whether approaches beyond ranking loss can also alleviate the gradient vanishing of negative samples and hence improve classification performance. We first study the following two methods: Focal Loss (Lin et al., 2017) and Negative Sampling (Mikolov et al., 2013). Then, we designed a novel approach called Combined-Contrastive to validate our perspective further.
4.5.1. Focal Loss
@@
It assigns higher weights to poorly classified samples (Lin et al., 2017), i.e., negative samples suffering from gradient vanishing in our scenario. Specifically, Focal Loss introduces a weight with hyper-parameter $\gamma$ to control the weight of samples:
(16) $\mathcal{L}_{\text{Focal}}=-\frac{1}{N}\sum_{i=1}^{N}[y_{i}(1-\hat{p}_{i})^{\gamma}\log(\hat{p}_{i})+(1-y_{i})\hat{p}_{i}^{\gamma}\log(1-\hat{p}_{i}))],$
@@
where $\gamma$ controls the relative weight. For those negative samples that may suffer from gradient vanishing when only using the BCE loss, their prediction score $\hat{p}_{i}$ should be wrongly high, making them have a higher weights $\hat{p}_{i}^{\gamma}$ than those negatives that are well-classified, i.e., with a low score. The larger $\gamma$, the higher the relative weights to those poorly-classified samples.
@@
We then conducted a set of experiments by comparing gradients and performance for Focal Loss with different $\gamma$. As shown in Fig. 7 (left), similar to Combined-Pair, Focal Loss also gets higher gradients on negative samples than the BCE loss. The larger the $\gamma$, the more weights on the poorly classified samples, and the larger the performance lift than the BCE method. These results validate that Focal Loss can also mitigate the gradient vanishing of negative samples and, hence, improve classification performance.
Figure 7. Evaluation of Focal loss and negative sampling. Left: Gradient norm (boxplot) and AUC (solid line) along with different $\gamma$ for Focal Loss. Right: AUC (blue) and BCE loss (red) after isotonic regression with increasingly aggressive Negative Sampling.
@@
Please note that the original $\hat{p_{i}}^{\gamma}$ is always less than $1.0$, and the gradients of negative samples in Focal loss are constantly smaller than BCE loss. For a fair comparison, we introduce a slight modification to the original formulation by replacing $\hat{p_{i}}^{\gamma}$ with $[\hat{p_{i}}^{\gamma}+(1-\sum_{k\in N_{-}}\hat{p_{k}}^{\gamma}/N_{-})]$. This adjustment normalizes the average weights of the negative samples to $1.0$, aligning it with the BCE loss. Formally, the modified version is defined as:
(17) $\mathcal{L}_{\text{Focal'}}=-\frac{1}{N}$ $\sum\limits_{i=1}^{N}[y_{i}(1-\hat{p}_{i})^{\gamma}\log(\hat{p}_{i})$
$+$ $[\hat{p_{i}}^{\gamma}+(1-\sum_{k=1}^{N_{-}}\hat{p_{k}}^{\gamma}/N_{-})](1-y_{i})\log(1-\hat{p}_{i}))],$
4.5.2. Negative Sampling
@@
Another trial is to reduce the proportion of negative samples through negative sampling (Mikolov et al., 2013), thereby increasing the estimated CTR. This may consequently increase the gradient of negative samples because their gradients are proportional to the estimated CTR (Eq. 3.2.1). However, downsampling negative samples may lead to information degradation, especially in datasets like Criteo, which has already undergone negative sampling. We analyze the AUC and BCE loss of negative sampling, and for a fair comparison, we report the calibrated results by isotonic regression (Chakravarti, 1989; Leeuw et al., 2009). We observe that as negative sampling becomes more aggressive, both the model’s ranking and classification capabilities deteriorate in Fig. 7 (right), which indicates that negative sampling fails to improve performance.
4.5.3. Combined-Contrastive: A New Method
Besides validating our perspective on existing methods, such as Combined-Pair and Focal Loss, we’d like to derive new methods based on our perspective. We speculate that introducing an auxiliary loss that considers the label information, i.e., an auxiliary supervised loss, may provide larger gradients than the Binary Cross-Entropy (BCE) loss itself and hence mitigate the gradient vanishing issue. To this end, inspired by the supervised contrastive learning (Khosla et al., 2020), we have devised a novel approach termed Combined-Contrastive Loss, which integrates the BCE loss with Contrastive Loss. Specifically, the Contrastive Loss is employed to encourage embeddings belonging to the same class to be closely grouped while ensuring distinct separation between embeddings from different classes. Formally,
(18) $\mathcal{L}^{\text{CC}}=\alpha\mathcal{L}_{\text{BCE}}+(1-\alpha)\mathcal{L}_{\text{Contr}},$
(19) $\mathcal{L}_{\text{Contr}}=\frac{1}{|N|}\sum_{i=1}^{N}\frac{-1}{|P(i)|}\sum_{p\in P(i)}\log\frac{\exp(\mathbf{z}_{i}\mathbf{z}_{p}/\tau)}{\sum\limits_{a\in A(i)}\exp(\mathbf{z}_{i}\mathbf{z}_{a}/\tau)},$
where $\mathcal{L}_{\text{CC}}$ and $\mathcal{L}_{\text{Contr}}$ represent the Combined-Contrastive Loss and the Contrastive Loss, respectively. $N$ denotes the number of samples in the batch, $A(i)$ denotes the whole samples set except $i$-th sample itself, $P(i)$ denotes a sample subset that contains all samples in $A(i)$ that are with the same label as $i$-th sample, $\mathbf{z_{i}}$ denotes the embedding of the $i$-th sample. We set $\alpha=0.9$ and $\tau=0.4$.
Table 2. Performance and gradient norm of negative samples of BCE method and Combined-Contrastive. The training stage’s average values on the first epoch are reported.
Stage Metrics BCE Method Combined-Contrastive
Training Gradient Norm $4.9\times 10^{-6}$ $\mathbf{7.5\times 10^{-6}}$
BCE loss $\downarrow$ 0.09667 0.09428
[…]
BCE $\downarrow$ 0.12152 0.12147
We conduct experiments on the same artificial Criteo dataset with $\beta_{\text{pos}}=0.1$ as mentioned in Sec. 4.1. As shown in Tab. 2, Combined-Contrastive gets a higher AUC and smaller BCE loss in the testing set, indicating better classification and ranking ability. In addition, similar to the Combined-Pair, it also gets lower training BCE loss and larger gradients than the BCE method. This verifies that by introducing the auxiliary contrastive loss, Combined-Contrastive can also mitigate the gradient vanishing issue and hence improve the classification ability.
@@ 5. Stability and Compatibility
Recent studies(Bai et al., 2023; Yan et al., 2022) have raised concerns regarding tailoring ranking loss into CTR prediction. Some studies (Yan et al., 2022) are concerned that ranking loss introduces score drifting, leading to optimization instability, while others (Bai et al., 2023) express concerns about compatibility issues between ranking loss and BCE loss. Both issues can also lead to negative effects on optimization. Hence, we aim to investigate whether classification-ranking loss, especially Combined-Pair, also has these two issues or not.
@@ 5.1. Stablility
Stability (Yan et al., 2022) refers to the phenomenon that the model scores may keep drifting during model training. Intuitively, singly employing ranking loss causes the model to focus solely on the relative orders between samples, neglecting the absolute prediction scores. This leads to score drifting (Yan et al., 2022), resulting in non-scale calibrated outcomes and exacerbating bias distribution. Hence, we wonder whether combining ranking loss with BCE loss in Combined-Pair may worsen the bias compared to using only BCE loss (i.e., the BCE method). It is worth mentioning that we did not employ any post-processing calibration techniques here. Otherwise, the bias would not reflect its training stability.
@@
We bucketize the samples with equal frequency of positive samples and plot the bias of corresponding samples within each bucket. As shown in Fig. 8, in both online and offline advertising, the BCE method severely underestimates the click-through rate (CTR) in the lower buckets, while Combined-Pair has a much smaller bias. In the higher buckets, the BCE method overestimates the CTR, and Combined-Pair also has a small bias in those buckets. The reason is probably that Combined-Pair mitigates the gradient vanishing issue, thus facilitating model optimization and leading to lower bias. In summary, Combined-Pair demonstrates superior calibration ability compared to the BCE method, avoiding the issue of score drifting and ensuring stability.
@@ 5.2. Compatibility
Recent research (Bai et al., 2023) argued that the pointwise and pairwise loss may not be compatible since they have different global minima. Thus, we are curious about the compatibility of the two losses within Combined-Pair: the BCE loss vs. the RankNet loss. We refrain from comparing their global minima (Bai et al., 2023), as differences in global minima among losses do not inherently indicate incompatibility. For instance, while l1 or l2 regularization losses do not share the same global minimum as most optimization objectives, they are still effectively utilized in tandem to mitigate overfitting.
@@
Negative $\frac{\hat{p}_{j}}{N}>0$ $\frac{\sum_{i=1}^{N_{+}}\sigma(z_{j}^{(-)}-z_{i}^{(+)})}{N_{+}N_{-}}>0$ Same
Positive $-\frac{1-\hat{p}_{i}}{N}<0$ $-\frac{\sum_{j=1}^{N_{-}}[1-\sigma(z_{i}^{(+)}-z_{j}^{(-)})]}{N_{+}N_{-}}<0$ Same
Instead, we investigate the compatibility of these losses by analyzing their gradients, examining whether there is a conflict in gradients between them. As shown in Tab. 3, the BCE loss and RankNet loss gradients are always aligned in the same sign, leading to the same optimization directions by two objectives. Therefore, the optimization directions of the combined-pair loss remain compatible without conflict.
@@ 6. Online Deployment
We conduct an empirical evaluation of Combined-Pair on the Click-Through Rate (CTR) prediction task across three distinct online advertising scenarios in Tencent: WeChat Channels, WeChat Moments, and the Demand-Side Platform (DSP).
6.1. Deployment Details
@@
The underlying model architecture utilizes Heterogeneous Experts with Multi-Embedding framework (Pan et al., 2024). Specifically, our approach involves training multiple feature interaction experts, such as GwPFM (Pan et al., 2024) (a variant of FFM (Juan et al., 2017) and FwFM (Pan et al., 2018a)), IPNN (Qu et al., 2016), DCN V2 (Wang et al., 2021), or FlatDNN, to capture diverse feature interactions for sparse ID features. Additionally, we learn multiple embedding tables for all features (Guo et al., 2024; Su et al., 2024; Pan et al., 2018b), with each table corresponding to one or several experts. For sequence features, we employ the TIN (Zhou et al., 2024) to capture the semantic-temporal correlation.
@@
Based on the above backbone architecture, we deployed Combined-Pair and conducted online A/B testing from early July 2023 to August 2023. We configured the Combined-Pair with $\alpha=0.9$ and adopted streaming training. The CTR varies from $0.1\%$ to $2.0\%$ in different scenarios.
The ranking loss adds an extra computation cost of $O(N_{+}\times N_{-})$ per batch, where $N_{+}$ and $N_{-}$ denote the number of positive and negative samples within a batch. Such additional computation cost is negligible compared with the complexity of the backbone model. Besides, the ranking loss only influences the training time and doesn’t affect the inference time. During the online A/B test, the inference time and QPS are consistent with the baseline.
6.2. Overall Performance
We examined the gradient distribution of negative samples and normalized their frequency, as shown in Fig. 9. The analysis reveals that the gradient distribution of negative samples for the Combined-Pair method is significantly right-skewed compared to the BCE method, indicating that the Combined-Pair method obtains larger gradients.
@@ 6.2. Overall Performance
Upon further examination of the model’s online performance metrics, we observed that the Combined-Pair method significantly improves all business metrics compared to the BCE method, as shown in Tab. 4. For instance, during one month of A/B testing with 20% traffic, the Combined-Pair method achieves a cost lift of 0.59% and a Gross Merchandise Value (GMV) lift of 0.70% over the BCE method in the WeChat Moments scenario.
In addition, we also examined the BCE loss for online deployment. Our observations indicate that the reduction in BCE loss for the Combined-Pair method ranges from 0.01% to 0.1% over a span of 7 days compared to the BCE method during the A/B testing period.
6.3. New Ads Performance
Given that new ads have only a few training samples and are more prone to optimization difficulties, we specifically focused on the performance of the Combined-Pair method in the WeChat Channels scenario. The results, as shown in Tab. 5, demonstrate a significant performance improvement achieved by the Combined-Pair method. Specifically, it achieves a GMV lift of 1.26% and a cost lift of 0.34%, which are statistically significant, as confirmed by the t-test.
@@ 7. Conclusion
In this paper, we have identified a challenge associated with using only binary-cross entropy loss for Click-Through Rate (CTR) prediction when positive feedback is sparse. Specifically, we have observed the issue of gradient vanishing for negative samples in such scenarios. To address this challenge, we propose a novel perspective by introducing an auxiliary ranking loss. We explain that the inclusion of this additional ranking loss leads to the generation of larger gradients for negative samples, effectively mitigating the problem of gradient vanishing. Through comprehensive experiments and analysis, we have provided strong evidence to support our perspective.
Acknowledgements.
We want to express our sincere gratitude to the following individuals for their invaluable contributions to this research regarding the analysis, online implementation, and experiments: Ming Yue, Kaixin Li, Zhixiang Feng, Xian Hu, Yaqian Zhang, Shifeng Wen, Jiancheng Wang, Yiding Deng, Zeen Xu, Xiaochen Wang, Chen Cai, GenBao Chen, Chaonan Guo, Junjie Zhai.